please help me solve this problem 20-4X=20-12
Answer by checkley71(8403)

20-4X=20-12 OR -4X=20-12-20 OR -4X=-12 OR X=-12/-4 OR X=3
PROOF 20-4*3=20-12 OR 20-12=20-12 OR 8=8